# BinaryAI: Binary Software Composition Analysis via Intelligent Binary Source Code Matching

### v2

@@ Abstract.
While third-party libraries (TPLs) are extensively reused to enhance productivity during software development, they can also introduce potential security risks such as vulnerability propagation. Software composition analysis (SCA), proposed to identify reused TPLs for reducing such risks, has become an essential procedure within modern DevSecOps. As one of the mainstream SCA techniques, binary-to-source SCA identifies the third-party source projects contained in binary files via binary source code matching, which is a major challenge in reverse engineering since binary and source code exhibit substantial disparities after compilation. The existing binary-to-source SCA techniques leverage basic syntactic features that suffer from redundancy and lack robustness in the large-scale TPL dataset, leading to inevitable false positives and compromised recall. To mitigate these limitations, we introduce BinaryAI, a novel binary-to-source SCA technique with two-phase binary source code matching to capture both syntactic and semantic code features. First, BinaryAI trains a transformer-based model to produce function-level embeddings and obtain similar source functions for each binary function accordingly. Then by applying the link-time locality to facilitate function matching, BinaryAI detects the reused TPLs based on the ratio of matched source functions. Our experimental results demonstrate the superior performance of BinaryAI in terms of binary source code matching and the downstream SCA task. Specifically, our embedding model outperforms the state-of-the-art model CodeCMR, i.e., achieving 22.54% recall@1 and 0.34 MRR compared with 10.75% and 0.17 respectively. Additionally, BinaryAI outperforms all existing binary-to-source SCA tools in TPL detection, increasing the precision from 73.36% to 85.84% and recall from 59.81% to 64.98% compared with the well-recognized commercial SCA product Black Duck.
@@ 1. Introduction
Software composition analysis (SCA) (Jiang et al., 2023a; Yuan et al., 2019; Synopsys, 2023) refers to identifying the open-source components (reused third-party libraries, i.e., TPLs) contained in the software artifacts for cost-effective development. Based on the SCA result, developers can easily track the security threats introduced to the software artifact by TPLs, such as vulnerability propagation and license violation (Synopsys, [n. d.]; Julie Peterson, 2021; Amrita Pathak, 2022; Liu et al., 2022). Considering diverse forms of target software project and identified TPLs, the existing SCA techniques are divided into multiple categories (e.g., binary-to-binary SCA (Tang et al., 2020; Yang et al., 2022; Tang et al., 2022a) and binary-to-source SCA (Duan et al., 2017; Yuan et al., 2019; Miyani et al., 2017)). In particular, as one of the mainstream SCA techniques, binary-to-source SCA techniques identify the source code projects as reused TPLs contained in the target binary file by measuring the similarity between the binary file and a large-scale collected TPL dataset based on their extracted code features (i.e., binary source code matching). During the modern DevSecOps CI/CD pipeline, binary-to-source SCA is normally integrated into the build or deploy phase to automatically scan the components along with security risks within binary files. Typically, binary-to-source SCA tends to be scalable and practical in real-world software development scenarios (Synopsys, 2023; Scantist, 2023; Red Hat, 2022) by conveniently incorporating new open source repositories into the TPL dataset.
@@
In this paper, we evaluate the effectiveness of BinaryAI in terms of binary source code matching and TPL detection (i.e., SCA). Specifically, we first construct three datasets: 1) training set for the model containing around 10M function pairs as positive samples, 2) large-scale TPL dataset to construct the SCA database and corpus for retrieving similar source functions, and 3) SCA test set with manually labeled components and binary-to-source function mappings. The evaluation results reveal that the binary source code matching model in BinaryAI outperforms the state-of-the-art model CodeCMR by increasing recall@1 from 11.92% to 22.73% for the SCA test set. Moreover, the locality-driven matching can effectively identify the correct source function from top-k retrieved results, further increasing recall@1 from 22.73% to 66.90% that is close to the upper bound 70.45% (i.e., recall@100) restricted by the model capability. Based on the matched source functions, we evaluate the accuracy of BinaryAI regarding TPL detection. The evaluation results demonstrate that BinaryAI dominates the performance among all the existing binary-to-source SCA tools, e.g., outperforming the start-of-the-art academic SCA tool B2SFinder by increasing the precision from 31.78% to 85.84% and the recall from 54.93% to 64.98%. It even outperforms the well-recognized commercial SCA product Black Duck by increasing the precision from 73.36% to 85.84% and the recall from 59.81% to 64.98%.
@@ 2.1. Software Composition Analysis
Software composition analysis (SCA) typically refers to identifying third-party libraries (TPLs) in the target software project to track security threats and license violations introduced by these open-source components. Given the potential risks to the software supply chain associated with accessing the source code (e.g., privacy policy), binary SCA has emerged as the predominant technique, which can be easily integrated into the build or deploy phase during DevSecOps to automatically scan the components in binary files (Duan et al., 2017; Yuan et al., 2019; Synopsys, 2023; Scantist, 2023). Existing binary SCA techniques (Duan et al., 2017; Yuan et al., 2019; Tang et al., 2020; Li et al., 2017b; Yang et al., 2022) extract software features from a large-scale TPL dataset to construct the SCA database and then utilize code clone detection to identify similar features between TPLs and the binary file. Subsequently, they recognize the TPLs as the reused components if the number of similar features exceeds a pre-defined threshold. Based on different forms of TPLs in the database, binary SCA can be divided into two categories: binary-to-source SCA (Duan et al., 2017; Yuan et al., 2019; Synopsys, 2023; Scantist, 2023) and binary-to-binary SCA (Tang et al., 2020, 2022a; Yang et al., 2022).
@@ 3.3. Locality-driven Matching
Input: bin2src_topk   normal-$\triangleright$ Retrieved topk similar source functions
@@
Result: bin2src_match   normal-$\triangleright$ Matched binary source function pairs
@@ 3 Function MatchFuncPairs:
7                   file2pairs[file].add(bin_rva normal-$\Rightarrow$ src_func) for file in src_files
@@ 4.2. Experiment Setup
For locality-driven matching, we evaluate the accuracy by identifying positive binary source function pairs, as well as its contribution to refining the top-1 result retrieved by the model, where we adopt the 15 binary files with manually labeled function mappings as the ground truth. Eventually, we utilize all the 150 binary files with 1,045 labeled reused TPLs for the SCA task. We compare BinaryAI with the existing binary-to-source SCA tools including two academic tools: OSSPolice (Duan et al., 2017) and B2SFinder (Yuan et al., 2019), and two well-established commercial products: Scantist (Scantist, 2023) and Synopsys Black Duck (Synopsys, 2023). For a fair comparison, we utilize the same TPL dataset (12,013 projects) to build their corresponding SCA database.
Note that we adopt the same metrics for investigating these two tasks that include Precision (i.e., the ratio of true positives to all the derived results), Recall (i.e., the ratio of true positives to all the ground-truth data), and F1 score (i.e., the measure of accuracy by considering both precision and recall). Considering the trade-off between precision and recall, we set the threshold $\theta$ to 0.01 that achieves the maximum F1 score [^3] We evaluate the impact of multiple reasonable threshold setups. Note that applying these thresholds does not incur significant performance variations, indicating the effectiveness of TPL detection for BinaryAI..
@@ 4.3.3. RQ3: Accuracy of TPL Detection
Lastly, we compare the performance of BinaryAI with the existing tools in terms of binary-to-source SCA. Table 3 demonstrates the overall result of TPL detection for 1,045 labeled components within 150 binary files. We can observe that, in general, BinaryAI significantly outperforms all the other SCA tools. For instance, BinaryAI significantly outperforms typical academic binary-to-source SCA techniques OSSPolice and B2SFinder. Furthermore, BinaryAI can even outperform well-recognized commercial binary SCA product Black Duck (85.84% vs. 73.36% precision, 64.98% vs. 59.81% recall, and 73.97% vs. 65.90% F1). Figure 6 presents the distribution of precision and recall for TPL detection across the 150 binary files. We observe that BinaryAI can dominate the precision and recall of component identification in most binary files, followed by the Black Duck. On the contrary, OSSPolice, B2SFinder, and Scantist cannot generalize well to our SCA test cases with compromised precision or recall.
@@
Tool Verification of TPL Detection #TP #FP #FN P (%) R (%) F1 (%) OSSPolice 348 191 697 64.56 33.30 43.94 B2SFinder 574 1232 471 31.78 54.93 40.26 Scantist 232 108 813 68.24 22.20 33.50 Black Duck 625 227 420 73.36 59.81 65.90 BinaryAI 679 112 366 85.84 64.98 73.97
@@ 5. Threats to Validity
Threats to external validity. The threat to external validity mainly lies in the subjects and dataset. To reduce this threat, we compare the model in BinaryAI with the state-of-the-art model CodeCMR in the domain of binary source code matching. For the downstream SCA task, we compare BinaryAI with four typical binary SCA tools from both industry (Scantist, Black Duck) and academia (OSSPolice, B2SFinder). For the dataset, we follow previous works (Woo et al., 2021; Tang et al., 2022b; Wang et al., 2022) to construct substantial binary source function pairs as the training dataset by building an automatic compilation pipeline for ArchLinux packages. Meanwhile, we collect a large-scale TPL dataset and create the largest corpus of source functions respectively in binary-to-source SCA and binary source code matching. Considering the lack of publicly available ground-truth data, it also takes the authors excessive manual effort to calibrate the ground-truth components and label the correspondence of functions.

### v3

@@ Abstract.
While third-party libraries (TPLs) are extensively reused to enhance productivity during software development, they can also introduce potential security risks such as vulnerability propagation. Software composition analysis (SCA), proposed to identify reused TPLs for reducing such risks, has become an essential procedure within modern DevSecOps. As one of the mainstream SCA techniques, binary-to-source SCA identifies the third-party source projects contained in binary files via binary source code matching, which is a major challenge in reverse engineering since binary and source code exhibit substantial disparities after compilation. The existing binary-to-source SCA techniques leverage basic syntactic features that suffer from redundancy and lack robustness in the large-scale TPL dataset, leading to inevitable false positives and compromised recall. To mitigate these limitations, we introduce BinaryAI, a novel binary-to-source SCA technique with two-phase binary source code matching to capture both syntactic and semantic code features. First, BinaryAI trains a transformer-based model to produce function-level embeddings and obtain similar source functions for each binary function accordingly. Then by applying the link-time locality to facilitate function matching, BinaryAI detects the reused TPLs based on the ratio of matched source functions. Our experimental results demonstrate the superior performance of BinaryAI in terms of binary source code matching and the downstream SCA task. Specifically, our embedding model outperforms the state-of-the-art model CodeCMR, i.e., achieving 22.54% recall@1 and 0.34 MRR compared with 10.75% and 0.17 respectively. Additionally, BinaryAI outperforms all existing binary-to-source SCA tools in TPL detection, increasing the precision from 73.36% to 85.84% and recall from 59.81% to 64.98% compared with the well-recognized commercial SCA product.
@@ 1. Introduction
Software composition analysis (SCA) (Jiang et al., 2023a; Yuan et al., 2019; Synopsys, 2023) refers to identifying the open-source components (reused third-party libraries, i.e., TPLs) contained in the software artifacts for cost-effective development. Based on the SCA result, developers can easily track the security threats introduced to the software artifact by TPLs, such as vulnerability propagation and license violation (Synopsys, [n. d.]; Julie Peterson, 2021; Amrita Pathak, 2022; Liu et al., 2022). Considering diverse forms of target software project and identified TPLs, the existing SCA techniques are divided into multiple categories (e.g., binary-to-binary SCA (Tang et al., 2020; Yang et al., 2022; Tang et al., 2022a) and binary-to-source SCA (Duan et al., 2017; Yuan et al., 2019; Miyani et al., 2017)). In particular, as one of the mainstream SCA techniques, binary-to-source SCA techniques identify the source code projects as reused TPLs contained in the target binary file by measuring the similarity between the binary file and a large-scale collected TPL dataset based on their extracted code features (i.e., binary source code matching). During the modern DevSecOps CI/CD pipeline, binary-to-source SCA is normally integrated into the build or deploy phase to automatically scan the components along with security risks within binary files. Typically, binary-to-source SCA tends to be scalable and practical in real-world software development scenarios (Synopsys, 2023; Duan et al., 2017; Red Hat, 2022) by conveniently incorporating new open source repositories into the TPL dataset.
@@
In this paper, we evaluate the effectiveness of BinaryAI in terms of binary source code matching and TPL detection (i.e., SCA). Specifically, we first construct three datasets: 1) training set for the model containing around 10M function pairs as positive samples, 2) large-scale TPL dataset to construct the SCA database and corpus for retrieving similar source functions, and 3) SCA test set with manually labeled components and binary-to-source function mappings. The evaluation results reveal that the binary source code matching model in BinaryAI outperforms the state-of-the-art model CodeCMR by increasing recall@1 from 11.92% to 22.73% for the SCA test set. Moreover, the locality-driven matching can effectively identify the correct source function from top-k retrieved results, further increasing recall@1 from 22.73% to 66.90% that is close to the upper bound 70.45% (i.e., recall@100) restricted by the model capability. Based on the matched source functions, we evaluate the accuracy of BinaryAI regarding TPL detection. The evaluation results demonstrate that BinaryAI dominates the performance among all the existing binary-to-source SCA tools, e.g., outperforming the start-of-the-art academic SCA tool B2SFinder by increasing the precision from 31.78% to 85.84% and the recall from 54.93% to 64.98%. It even outperforms the well-recognized commercial SCA product by increasing the precision from 73.36% to 85.84% and the recall from 59.81% to 64.98%.
@@ 2.1. Software Composition Analysis
Software composition analysis (SCA) typically refers to identifying third-party libraries (TPLs) in the target software project to track security threats and license violations introduced by these open-source components. Given the potential risks to the software supply chain associated with accessing the source code (e.g., privacy policy), binary SCA has emerged as the predominant technique, which can be easily integrated into the build or deploy phase during DevSecOps to automatically scan the components in binary files (Duan et al., 2017; Yuan et al., 2019; Synopsys, 2023). Existing binary SCA techniques (Duan et al., 2017; Yuan et al., 2019; Tang et al., 2020; Li et al., 2017b; Yang et al., 2022) extract software features from a large-scale TPL dataset to construct the SCA database and then utilize code clone detection to identify similar features between TPLs and the binary file. Subsequently, they recognize the TPLs as the reused components if the number of similar features exceeds a pre-defined threshold. Based on different forms of TPLs in the database, binary SCA can be divided into two categories: binary-to-source SCA (Duan et al., 2017; Yuan et al., 2019; Synopsys, 2023) and binary-to-binary SCA (Tang et al., 2020, 2022a; Yang et al., 2022).
@@ 3.3. Locality-driven Matching
Input: bin2src_topk
  $\triangleright$ Retrieved topk similar source functions
@@
Result: bin2src_match
  $\triangleright$ Matched binary source function pairs
@@ 4 Function MatchFuncPairs:
8                   file2pairs[file].add(bin_rva $\Rightarrow$ src_func) for file in src_files
@@ 4.2. Experiment Setup
For locality-driven matching, we evaluate the accuracy by identifying positive binary source function pairs, as well as its contribution to refining the top-1 result retrieved by the model, where we adopt the 15 binary files with manually labeled function mappings as the ground truth. Eventually, we utilize all the 150 binary files with 1,045 labeled reused TPLs for the SCA task. We compare BinaryAI with the existing binary-to-source SCA tools including two academic tools: OSSPolice (Duan et al., 2017) and B2SFinder (Yuan et al., 2019), and two well-established commercial products: bsca-A and bsca-B [^3] The specific commercial tools are not disclosed due to the constraints of confidentiality agreements and the proprietary nature of the software, thus the names of the commercial tools used in this paper have been anonymized.. For a fair comparison, we utilize the same TPL dataset (12,013 projects) to build their corresponding SCA database.
Note that we adopt the same metrics for investigating these two tasks that include Precision (i.e., the ratio of true positives to all the derived results), Recall (i.e., the ratio of true positives to all the ground-truth data), and F1 score (i.e., the measure of accuracy by considering both precision and recall). Considering the trade-off between precision and recall, we set the threshold $\theta$ to 0.01 that achieves the maximum F1 score.
@@ 4.3.3. RQ3: Accuracy of TPL Detection
Lastly, we compare the performance of BinaryAI with the existing tools in terms of binary-to-source SCA. Table 3 demonstrates the overall result of TPL detection for 1,045 labeled components within 150 binary files. We can observe that, in general, BinaryAI significantly outperforms all the other SCA tools. For instance, BinaryAI significantly outperforms typical academic binary-to-source SCA techniques OSSPolice and B2SFinder. Furthermore, BinaryAI can even outperform well-recognized commercial binary SCA product bsca-B (85.84% vs. 73.36% precision, 64.98% vs. 59.81% recall, and 73.97% vs. 65.90% F1). Figure 6 presents the distribution of precision and recall for TPL detection across the 150 binary files. We observe that BinaryAI can dominate the precision and recall of component identification in most binary files, followed by the bsca-B. On the contrary, OSSPolice, B2SFinder, and bsca-A cannot generalize well to our SCA test cases with compromised precision or recall.
@@
Tool Verification of TPL Detection #TP #FP #FN P (%) R (%) F1 (%) OSSPolice 348 191 697 64.56 33.30 43.94 B2SFinder 574 1232 471 31.78 54.93 40.26 bsca-A 232 108 813 68.24 22.20 33.50 bsca-B 625 227 420 73.36 59.81 65.90 BinaryAI 679 112 366 85.84 64.98 73.97
@@ 5. Threats to Validity
Threats to external validity. The threat to external validity mainly lies in the subjects and dataset. To reduce this threat, we compare the model in BinaryAI with the state-of-the-art model CodeCMR in the domain of binary source code matching. For the downstream SCA task, we compare BinaryAI with four typical binary SCA tools from both industry and academia (OSSPolice, B2SFinder). For the dataset, we follow previous works (Woo et al., 2021; Tang et al., 2022b; Wang et al., 2022) to construct substantial binary source function pairs as the training dataset by building an automatic compilation pipeline for ArchLinux packages. Meanwhile, we collect a large-scale TPL dataset and create the largest corpus of source functions respectively in binary-to-source SCA and binary source code matching. Considering the lack of publicly available ground-truth data, it also takes the authors excessive manual effort to calibrate the ground-truth components and label the correspondence of functions.